### v1

@@ Abstract
We present Hapax Locks, a novel locking algorithm that is simple, enjoys constant-time arrival and unlock paths, provides FIFO admission order, and which is also space efficient and generates relatively little coherence traffic under contention in the common case. Hapax Locks offer performance (both latency and scalability) that is comparable with the best state of the art locks, while at the same time Hapax Locks impose fewer constraints and dependencies on the ambient runtime environment, making them particularly easy to integrate or retrofit into existing systems or under existing application programming interfaces Of particular note, no pointers shift or escape ownership between threads in our algorithm. 111Extended version: https://arxiv.org/abs/0000.00000.
@@
††margin: Context; Motivation; background; frame the problem; ††margin: endow; imbueLocks often have a crucial impact on the performance of parallel software, hence they remain the focus of intensive research with a steady stream of algorithms proposed over the last several decades. Picking an algorithm involves striking a balance between a number of factors, such as : latency in the uncontended case; throughput in the contended case; admission fairness; space requirements; coherence traffic induced by the lock; any dependencies the lock might have on other subsystems (such as dynamic memory allocation); and how well an algorithm might be retrofit under existing locking application interfaces. Our investigation focuses on algorithms and implementations thereof that gracefully tolerate large numbers of dynamically created and destroyed threads and locks, and that can be implemented easily under C++ locking interfaces and POSIX pthreads APIs. Applications expect to be able to acquire and hold multiple locks and to release in any order. In addition, we prefer to avoid dynamic memory allocation, and the requirement for explicit constructor and destructor methods for threads and lock instances. In this paper, we present Hapax Locks, which possess all aforementioned properties.
@@ 2 Tidex Lock Algorithm
Our design starts with and is inspired by the Tidex[ppopp17-Ramalhete] mutual exclusion algorithm. Briefly, in Tidex, each lock instance has Arrive and Depart fields which contain thread identifier values. The lock is in unlocked state if and only if Arrive equals Depart. On arrival in the lock method, a thread atomically exchanges its own unique identity value into Arrive, which returns the predecessor value. The thread then waits for Depart to become equal to that predecessor value that it has in hand, at which point it has become the owner and can enter the critical section. In the corresponding unlock operator, the thread just stores its own thread identity value into Depart. That is, a thread in unlock announces that it has surrendered ownership by posting its own identity value into Depart, and that outgoing thread nevers knows the identity of its successor, if any exists. Intuitively, when a thread is waiting and its predecessor departs and deposits its identity into Depart, that thread becomes the next owner. Similar to CLH[craig-clh], the queue of waiting threads is implicit, with no manifest or explicit in-memory linked list. Waiting threads, if any, know only the identity of their immediate predecessor and wait for that value to appear in the Depart field. Collectively, however, a queue is formed. Identifier values are expected to be temporally unique and a given value can be associated with at most one thread at any given time.
@@ 3 Hapax Locks
Hapax Locks are based on, but address the aforementioned performance shortcomings inherent in Tidex Locks. With Hapax Locks, threads no longer bother to maintain their own ID and alternative ID values. In lock, our thread notionally executes an atomic fetch_add(1) against a global shared 64-bit ID generator to conjure a new unique 64-bit value, and we use that as the thread ID for the acquire-release episode. The single generator instance is used for all threads and all locks in a process. The value from the generator is our hapax333In linguistics, hapax legomenon, – or just hapax (απαξ “once”) – refers to a word that appears just once in a body of text. In Hapax Locks, a particular hapax value is a nonce and appears at most once in the execution history of a program. value. Instead of a thread just having two Tidex-style identity values, it can now have multiple, allocated dynamically, one unique for each lock acquisition operation. This approach avoids any need to fetch and check Depart to ensure that our value does not happen to already appear in Arrive or Depart. A given hapax value, once allocated and made public by installation into Arrive, is never reused or recycled in the history of the process. With 64-bit values, we don’t, as a practical concern, need to worry about overflow and wrap-around of the hapax values generated. Unique hapax values are plentiful, and for all practical purposes, limitless.
@@
Both Tidex and Ticket locks use global spinning, which can result in potentially excessive write invalidation rates 444Imagine a contended Ticket Lock with 10 waiting threads, and no arriving threads. When the owner surrenders the lock, the write will invalidate, at worst case, in 10 remote caches. Only one of those waiting threads that suffered invalidation will become the next owner and the remainder will continue to busy-wait, pulling the line back into their caches after the invalidation. When that new owner ultimately surrenders ownership, its write will invalidate in 9 remote caches, and so on. So if we have waiting threads, each with a private multiple-reader-exclusive-writer MESI-style cache, then the number of invalidations required to pass ownership to all the threads, assuming no new arrivals, is at least – potentially quadratic write invalidation rates. We refer to the number of caches that require invalidation from a given store instruction as the invalidation set or invalidation blast zone of that store. Broadly, larger invalidation sets may incur more latency and consume more interconnect bandwidth.. In Tidex, for instance, all waiting threads for a given lock busy wait on the Depart field. To avoid such global spinning, we take inspiration from TWA[EuroPar19-TWA, arxiv-TWA] – itself derived from ticket locks – and implement a shared waiting array, where all threads and all locks use the same shared array. We opted to provision the array with 4096 slots. In theory, the number of slots should be a function of the number of logical processors in the system, but 4096 is reasonable even on very large modern systems. Threads perform long-term waiting via elements in the array. TWA hashed the ticket value and lock address to form an index into the waiting array, and the elements in those array slots were slot-specific sequence numbers. When TWA waiters observe the slot sequence number change, they would then recheck the actual ticket lock Grant “now serving” field to see if they were granted ownership. We refer to this as proxy waiting. Updates to the array elements constitute a conservative hint and notification to waiters that the actual location of interest may have been modified – the waiter is then expected to check the authoritative “ground truth” variable. In the TWA unlock operator, after updating Grant, TWA then atomically increments the corresponding slot sequence number to signal any waiters. Note, however, that because of collisions, waiters observing the sequence number change can not safely intuit that Grant had shifted to the waited-for value, so waiters are required to conservatively recheck Grant and, if necessary, resume the proxy waiting protocol. Absent hash collisions, we have private waiting, but as collisions may manifest, we say we have semi-private waiting.
@@
For Hapax Locks, we could just apply any general hash function, taking the hapax value and, optionally, the lock address, to form an index into the array. That would certainly suffice, but a better option is available. In the mapping function from hapax to waiting array index, we extract the bits in the hapax associated with the thread’s block and discard the rest, and apply a simple cache-aware lightweight hash to form the index. (We multiply by 17[WeylSequence] to map adjacent block zones onto different cache sectors or lines in the array, to reduce false sharing). So, for better performance, the hash — mapping function from waited-upon hapax value to waiting array index — is intentionally aware of the block allocation scheme. While not strictly required, we believe this is a useful optimization and strikes a balance between temporal cache locality (re-use by the same thread) in the array and the risk of collisions between threads. This approach is illustrated in the ToSlot function in Listing-LABEL:Listing:Hapax-Locks.
@@
Under contention, the value of the current outgoing owner’s hapax variable, executing in unlock, is the same as the incoming successor’s pred variable, as found in the successor’s lock phase. In Ticket locks, a thread knows that its immediate waiting successor must hold a ticket value of just one more than the ticket value allocated to the current owner. In Tidex and Hapax Locks, however, the departing thread has no particular knowledge about the successor, other than being able to determine if it exists.
@@ 4 Hapax Lock Variation with Visible Waiters
We say the baseline version of Hapax Locks shown in Listing-LABEL:Listing:Hapax-Locks uses invisible waiters as waiters only monitor the waiting array, but never write into the array, whereas in the version described in this section, with visible waiters, waiters explicit announce their existence by updating slots in the waiting array.
@@
In Listing-LABEL:Listing:Nonce-Visible-Waiters, accesses to shared globals in the contented fast path are annotated in purple and the key fast-path positive handover accesses are shown in blue. These reflect the dominant mode of execution under sustained contention. Critically, the atomic CAS (compare-and-swap) operation in unlock efficiently and expeditiously transfers ownership to its waiting successor running in lock.
@@
Recall that we ensure a hapax value of 0 is never allocated. Normally slots in the array are 0 indicating they are unoccupied. Threads arriving to wait first try to CAS the slot associated with the waited-upon value (which is non-0) from 0 to that value. If that CAS fails, because the slot is occupied by another waiter, they then revert to degenerate Tidex-style global spinning. This should be rare and only occur on hash collisions in the array. Otherwise, that waiting thread owns and has reserved the slot for the duration of its waiting phase and then busy waits for the value to change to any other value. The corresponding unlock operator tries to clear the slot back to 0 with CAS. The slot value can then flicker to 0 and then to other values, because of concurrent activities, so the reader (waiter) is safe to interpret any change in the value as indicating that ownership has been conveyed. This protocol is safe by virtue of the hapax value non-recurring property. The slot can not ever revert to a previous value, so we do not miss wakeups to a waiter.
@@
This variation effects contended lock ownership handover via 2 distinct modes : (a) The outgoing thread, executing in the unlock phase, uses an atomic compare-and-swap (CAS), on the associated slot in the array to attempt to overwrite its own hapax value with 0. The particular slot in the array is determined via a hash on the hapax value. Recall that the value of the current outgoing owner’s hapax variable, in unlock, is the same as the incoming successor’s pred variable in the successor’s lock phase. If the CAS is successful, the outgoing thread has accomplished expedited handover and foregoes updating the Depart field. The handoff is assured (positive) as, by virtue of the CAS, the departing thread knows with certainty that the waiter exists, and is waiting at the slot, and will be able to observe the change, providing a synchronous rendezvous between the departing thread and that successor. The waiting successor detects positive handover by observing that the hapax value it installed in the array, shifts, by virtue of the CAS by the predecessor, and can avoid checking the Depart field, and can directly enter the critical section. Such expedited positive handover – accelerated departure – is the preferred mode of succession, enjoying the shortest paths and least latency. Critically, under sustained contention, neither the outgoing thread, running in unlock, nor the incoming successor, waiting in lock, access any locations in the shared lock body; Hapax Locks shares this beneficial attribute with MCS[tocs91-MellorCrummey] and CLH[craig-clh, ipss94-magnusson]. The outgoing thread, in unlock, does not need to store into Depart and the incoming successor, in lock, also does not need to fetch Depart to confirm handover, so we reduce coherence traffic on cache lines underlying the lock body. Unlock only needs to update the Depart field when there is no contention or if there were collisions in the waiting array. (b) The attempt at positive expedited handover fails – the CAS fails – so the outgoing thread reverts to the Tidex protocol and updates the Depart field, which is observed by the successor. Expedited positive handover can fail because of hash collisions on hapax values in the array, or if the unlock operation commences before the tardy successor has had a chance to register and make itself visible in the waiting array. The unlock operation can not determine with certainty if a successor exists – if the lock is contended or not – so it reverts to a more conservative protocol to release the lock.
@@
With positive handover, unlock can also skip updating the Depart (Egress) global if it can positively and definitively detect that it accomplished handover to a waiting successor. Furthermore, the successor can then also skip checking the Depart field. This optimization reduces write invalidations on the shared Depart variable. unlock only needs to update the Depart field when there is no contention or if there were collisions in the waiting array.
@@
The positive handover optimization enabled by visible waiters affords the most relative benefit under high flux situations, with high arrival and departure traffic rates on a given lock. If threads fail to acquire the lock during the brief initial visible spinning phase, then we know that throughput and progress over the contented lock is low, so the benefit of visible waiting and positive handover will not be as pronounced, so reverting to invisible waiting does not appreciably degrade performance in this particular operating regime. But by shifting to invisible waiting, our threads can reduce load on the visible waiting array, and allow other threads and locks to benefit from using positive handover via the primary array.
@@ 5 Related Work
CLH is extremely simple, has excellent Remote Memory Reference (RMR) complexity[stoc08-attiya, podc02-anderson], and requires just a single atomic exchange operation in the Acquire operation and no atomic read-modify-write instructions in Release. Unfortunately the waiting elements migrate between threads, which may be inimical to performance in NUMA environments. CLH locks also require explicit constructors and destructors, which may be inconvenient or preclude their use in some environments 555Many lock implementations require just trivial constructors to set the lock fields to 0 or some constant, and trivial destructors, which do nothing. The Linux kernel spinlock_t/qspinlock_t interface provides a constructor, but does not even expose a destructor. Such decisions influence which lock algorithms can be readily implemented under a particular interface.. Similarly, C++ std::mutex is allowed to be trivially destructible, meaning storage occupied by trivially destructible objects may be reused without calling the destructor. Under both GCC g++ version 15 and Clang++ Version 20, is_trivially_destructible reports True for std::mutex, and as such, destructors do not run.
@@
HemLock combines aspects of both CLH and MCS to form a lock that has very simple waiting node memory lifecycle constraints, is context-free but still scales well in common usage scenarios. HemLock requires just a singleton waiting element (queue node) per thread, which can be placed in thread-local storage. HemLock does not provide constant remote memory reference (RMR) complexity [opodis17-dvir] in scenarios where a thread holds multiple contended locks. In this situation the single node suffers from multi-waiting[spaa21-Dice]. Similar to MCS, HemLock lacks a constant-time unlock operation, whereas the unlock operator for CLH and Tickets is constant-time. Unlike MCS, HemLock requires active synchronous back-and-forth communication in the unlock path between the outgoing thread and its successor to protect the lifecycle of the waiting element. We note, however, that HemLock remains constant-time in the Release operator to the point where ownership is conveyed to the successor. HemLock uses address-based transfer of ownership, writing the address of the lock instead of a boolean, differentiating it from MCS and CLH. Reciprocating Locks[arxiv-recipro, ppopp25-dice], like HemLock, requires just a singleton per-thread waiting element allocated in thread-local storage. Reciprocating Locks has constant-time arrival and unlock paths, but is not FIFO, instead providing 2-bounded-bypass.
@@
Spinning : L indicates that waiters spin on a location that is usually collocated in the system topology with the waiting thread, which may confer performance benefits on some systems, such as those that use the Intel UPI[UPI] home-based coherence protocol, while U indicates that the location being waited upon is generally “unknown” and may be either local or remote to the waiting thread. P indicates private spinning, where at most one thread spins on a given location at a given time, G indicates global spinning, where all threads waiting on a given lock busy-wait on a single location, and S indicate semi-local[spaa21-Dice] spinning, where threads, absent hash collisions, enjoy private spinning.
@@
Context-free : indicates additional information does not need to be transferred from the lock operator to the corresponding Release operation.
@@
Explicit CTOR/DTOR Required : indicates the lock requires non-trivial constructors or destructors. CLH, for instance, requires destructors to run to release the wait elements referenced in the lock, to avoid memory leaks.
Thread CTOR/DTOR Required : indicates if per-thread constructors or destructors are needed to initialize or cleanup after dynamically created threads. We note that MCS implementations, could, if necessary, allocate queue nodes as needed, dynamically, in lock, and then release such queue nodes in unlock. This avoids the need for thread-level destructors. (To avoid circular or recursive dependencies, we further require a specialized node allocator that is implemented via different types of locks). To avoid the allocator becoming a contention hotspot and bottleneck, common MCS implementations use per-thread caches of available elements, which need to be released when the thread terminates. Similarly, CLH could, in theory, allocate and free queue nodes on demand, although most implementations use a thread-local cache containing a singleton pointer to a free queue node, as seen in Scott’s [ScottB24] Figure 4.14. MCSH[MCSH] and K42[K42, ScottB24] MCS variations side step this concern by allocating elements on-stack, but at the cost of added complexity.
Max Remote Misses per episode : is the worst-case maximum number of misses to remote memory incurred, under simple sustained contention, by a matching Acquire-Release pair. Misses to remote memory – memory homed on a different NUMA node – may be more expensive than local misses on various platforms, such as modern Intel systems that use the UPI coherence fabric [UPI], where miss requests are first adjudicated by the home node of the cache line. Algorithms where nodes circulate appear more vulnerable to accumulating such remote misses. For HemLock, we assume simple contention with no multi-waiting. We derived the “Max Remote Misses per episode” value via static inspection of the code, and determining which of the accesses that might cause coherence misses might also be to locations homed on remote nodes. We assume that back-to-back accesses of shared variables that reside on the same cache line, by a given thread, will result in just one miss. And for Hapax Locks we assume that the hapax value does not need to be replenished, which would incur an addition likely remote miss for the fetch-and-add atomic operation. We further assume no collisions (or near collisions, which could induce false sharing) in the waiting array.
@@
Dynamic Memory Allocation : indicates if the algorithm requires some type of dynamic memory allocation, typically for waiting elements.

### v2

@@ Abstract
We present Hapax Locks, a novel locking algorithm that is simple, enjoys constant-time arrival and unlock paths, provides FIFO admission order, and which is also space efficient and generates relatively little coherence traffic under contention in the common case. Hapax Locks offer performance (both latency and scalability) that is comparable with the best state of the art locks, while at the same time Hapax Locks impose fewer constraints and dependencies on the ambient runtime environment, making them particularly easy to integrate or retrofit into existing systems or under existing lock application programming interfaces Of particular note, no pointers shift or escape ownership between threads in our algorithm. 111source code:/https://github.com/davedice/Hapax-Locks.
@@
††margin: Context; Motivation; background; frame the problem; ††margin: endow; imbueLocks often have a crucial impact on the performance of parallel software, hence they remain the focus of intensive research with a steady stream of algorithms proposed over the last several decades. Picking an algorithm involves striking a balance between a number of factors, such as : latency in the uncontended case; throughput in the contended case; admission fairness; space requirements; coherence traffic induced by the lock; any dependencies the lock might have on other subsystems (such as dynamic memory allocation); and how well an algorithm might be retrofit under existing locking application interfaces. Our investigation focuses on algorithms and implementations thereof that gracefully tolerate large numbers of dynamically created and destroyed threads and locks, and that can be implemented easily under C++ locking interfaces and POSIX pthreads APIs. Applications expect to be able to acquire and hold multiple locks and to release in any order – scoped or balanced locking is not required and we have explicit lock and unlock operators. In addition, we prefer to avoid dynamic memory allocation, and the requirement for explicit constructor and destructor methods for threads and lock instances. In this paper, we present Hapax Locks, which possess all aforementioned properties.
@@ 2 Tidex Lock Algorithm
Our design starts with and is inspired by the Tidex[ppopp17-Ramalhete] mutual exclusion algorithm. Briefly, in Tidex, each lock instance has Arrive and Depart fields which contain thread identifier values. The lock is in unlocked state if and only if Arrive equals Depart. On arrival in the lock method, a thread atomically exchanges its own unique identity value into Arrive, which returns the predecessor value. The thread then waits for Depart to become equal to that predecessor value that it has in hand, at which point it has become the owner and can enter the critical section. In the corresponding unlock operator, the thread just stores its own thread identity value into Depart. That is, a thread in unlock announces that it has surrendered ownership by posting its own identity value into Depart – that outgoing thread never knows the identity of its successor, if any exists. Intuitively, when a thread is waiting and its predecessor departs and deposits its identity into Depart, that waiting thread becomes the next owner. Similar to CLH[craig-clh], the queue of waiting threads is implicit, with no manifest or explicit in-memory linked list. Waiting threads, if any, know only the identity of their immediate predecessor and wait for that value to appear in the Depart field. Collectively, however, a queue is formed. Identifier values are expected to be temporally unique and a given value can be associated with at most one thread at any given time.
@@ 3 Hapax Locks
Hapax Locks are based on, but address the aforementioned performance shortcomings inherent in Tidex Locks. With Hapax Locks, threads no longer bother to maintain their own ID and alternative ID values. In lock, our thread notionally executes an atomic fetch_add(1) against a global shared 64-bit ID generator to conjure a new unique 64-bit value, and we use that as the thread ID for the acquire-release episode. The single generator instance is used for all threads and all locks in a process. The value from the generator is our hapax333In linguistics, hapax legomenon, – or just hapax (απαξ “once”) – refers to a word that appears just once in a body of text. In Hapax Locks, a particular hapax value is a nonce and appears at most once in the execution history of a process. value. Instead of a thread just having two Tidex-style identity values, it can now have multiple, allocated dynamically, one unique for each lock acquisition operation. This approach avoids any need to fetch and check Depart to ensure that our value does not happen to already appear in Arrive or Depart. A given hapax value, once allocated and made public by installation into Arrive, is never reused or recycled in the history of the process. With 64-bit values, we don’t, as a practical concern, need to worry about overflow and wrap-around of the hapax values generated. Unique hapax values are plentiful, and for all practical purposes, limitless.
@@
Both Tidex and Ticket locks use global spinning, which can result in potentially excessive write invalidation rates 444Imagine a contended Ticket Lock with 10 waiting threads, and no arriving threads. When the owner surrenders the lock, the write will invalidate, at worst case, in 10 remote caches. Only one of those waiting threads that suffered invalidation will become the next owner and the remainder will continue to busy-wait, pulling the line back into their caches after the invalidation. When that new owner ultimately surrenders ownership, its write will invalidate in 9 remote caches, and so on. So if we have waiting threads, each with a private multiple-reader-exclusive-writer MESI-style cache, then the number of invalidations required to pass ownership to all the threads, assuming no new arrivals, is at least – potentially quadratic write invalidation rates. We refer to the number of caches that require invalidation from a given store instruction as the invalidation set or invalidation blast zone of that store. Broadly, larger invalidation sets may incur more latency and consume more interconnect bandwidth.. In Tidex, for instance, all waiting threads for a given lock busy wait on the Depart field. To avoid such global spinning, we take inspiration from TWA[EuroPar19-TWA, arxiv-TWA] – itself derived from ticket locks – and implement a shared waiting array, where all threads and all locks in a process use the same shared array. We opted to provision the array with 4096 slots. In theory, the number of slots should be a function of the number of logical processors in the system, but 4096 is reasonable even on very large modern systems. Threads perform long-term waiting via elements in the array. TWA hashed the ticket value and lock address to form an index into the waiting array, and the elements in those array slots were slot-specific sequence numbers. When TWA waiters observe the slot sequence number change, they would then recheck the actual ticket lock Grant “now serving” field to see if they were granted ownership. We refer to this as proxy waiting. Updates to the array elements constitute a conservative hint and notification to waiters that the actual location of interest may have been modified – the waiter is then expected to check the authoritative “ground truth” variable. In the TWA unlock operator, after updating Grant, TWA then atomically increments the corresponding slot sequence number to signal any waiters. Note, however, that because of collisions, waiters observing the sequence number change can not safely intuit that Grant had shifted to the waited-for value, so waiters are required to conservatively recheck Grant and, if necessary, resume the proxy waiting protocol. Absent hash collisions, we have private waiting, but as collisions may manifest, we say we have semi-private waiting.
@@
For Hapax Locks, we could just apply any general hash function, taking the hapax value and, optionally, the lock address, to form an index into the waiting array. That would certainly suffice, but a better option is available. In the mapping function from hapax to waiting array index, we extract the high-order bits in the hapax, which are associated with the thread, and discard the rest (the offset into the thread’s block), and then apply a simple cache-aware lightweight hash to form the index. (We multiply by 17[WeylSequence] to map adjacent block zones onto different cache sectors or lines in the array, to reduce false sharing). So, for better performance, the hash — mapping function from waited-upon hapax value to waiting array index — is intentionally aware of the block allocation scheme. While not strictly required, we believe this is a useful optimization and strikes a balance between temporal cache locality (re-use by the same thread) in the array and the risk of collisions or near collisions (false sharing) between threads. This approach is illustrated in the ToSlot function in Listing-LABEL:Listing:Hapax-Locks.
@@
Under contention, the value of the current outgoing owner’s hapax variable, executing in unlock, is the same as the incoming immediate successor’s pred variable, as found in the successor’s lock phase. In Ticket locks, a thread knows that its immediate waiting successor must hold a ticket value of just one more than the ticket value allocated to the current owner. In Tidex and Hapax Locks, however, the departing thread has no particular knowledge about the successor, other than being able to determine if it exists.
@@ 4 Hapax Lock Variation with Visible Waiters
We say the baseline version of Hapax Locks shown in Listing-LABEL:Listing:Hapax-Locks uses invisible waiters as waiters only monitor the waiting array, but never write into the array, whereas in the version described in this section, with visible waiters, waiters explicitly announce their existence by updating slots in the waiting array.
@@
In Listing-LABEL:Listing:Nonce-Visible-Waiters, accesses to shared globals in the contented fast path are annotated in purple and the key fast-path assured positive handover accesses are shown in blue. These reflect the dominant mode of execution under sustained contention. Critically, the atomic CAS (compare-and-swap) operation in unlock efficiently and expeditiously transfers ownership to its waiting successor running in lock.
@@
By convention, we ensure a hapax value of 0 is never allocated. Normally slots in the array are 0 indicating they are unoccupied. Threads arriving to wait first try to CAS the slot associated with the waited-upon value (which is non-0) from 0 to that value. If that CAS fails, because the slot is occupied by another waiter, they then revert to degenerate Tidex-style global spinning. This should be rare and only occur on hash collisions in the array. Otherwise, that waiting thread owns and has reserved the slot for the duration of its waiting phase and then busy waits for the value to change to any other value. The corresponding unlock operator tries to clear the slot back to 0 with CAS. The slot value can then flicker to 0 and then to other values, because of concurrent activities, so the reader (waiter) is safe to interpret any change in the value as indicating that ownership has been conveyed. This protocol is safe by virtue of the hapax value non-recurring property. The slot can not ever revert to a previous value, so we do not miss wakeups to a waiter.
@@
This variation effects contended lock ownership handover via 2 distinct modes : (a) The outgoing thread, executing in the unlock phase, uses an atomic compare-and-swap (CAS), on the associated slot in the array to attempt to overwrite its own hapax value with 0. The particular slot in the array is determined via a hash on the hapax value. Recall that the value of the current outgoing owner’s hapax variable, in unlock, is the same as the incoming successor’s pred variable in the successor’s lock phase. If the CAS is successful, the outgoing thread has accomplished expedited handover and foregoes updating the Depart field. The handoff is assured (positive) as, by virtue of the CAS, the departing thread knows with certainty that the waiter exists, and is waiting at the slot, and will be able to observe the change, providing a synchronous rendezvous between the departing thread and that successor. The waiting successor detects positive handover by observing that the hapax value it installed in the array, shifts, by virtue of the CAS by the predecessor, and can avoid checking the Depart field, and can directly enter the critical section. Such expedited positive handover – accelerated departure – is the preferred mode of succession, enjoying the shortest paths and least latency. Critically, under sustained contention, neither the outgoing thread, running in unlock, nor the incoming successor, waiting in lock, access any locations in the shared lock body; Hapax Locks shares this desirable attribute with MCS[tocs91-MellorCrummey] and CLH[craig-clh, ipss94-magnusson]. The outgoing thread, in unlock, does not need to store into Depart and the incoming successor, in lock, also does not need to fetch Depart to confirm handover, so we reduce coherence traffic on cache lines underlying the lock body. Unlock only needs to update the Depart field when there is no contention or if there were collisions in the waiting array. (b) The attempt at positive expedited handover fails – the CAS fails – so the outgoing thread reverts to the Tidex protocol and updates the Depart field, which is observed by the successor. Expedited positive handover can fail because of hash collisions on hapax values in the array, or if the unlock operation commences before the tardy successor has had a chance to register and make itself visible in the waiting array. The unlock operation can not determine with certainty if a successor exists – if the lock is contended or not – so it reverts to a more conservative protocol to release the lock.
@@
With positive handover, unlock can also skip updating the Depart field during egress. if it can positively and definitively detect that it accomplished handover to a waiting successor. Furthermore, the successor can then also skip checking the Depart field. This optimization reduces write invalidations on the shared Depart variable. unlock only needs to update the Depart field when there is no contention or if there were collisions in the waiting array.
@@
The positive handover optimization enabled by visible waiters affords the most relative benefit under high flux situations, with high arrival and departure traffic rates on a given lock. If threads fail to acquire the lock during the brief initial visible spinning phase, then we know that throughput and progress over the contented lock is low, so the benefit of visible waiting and positive handover will not be as pronounced, so reverting to invisible waiting does not appreciably degrade performance in this particular operating regime. But by shifting to invisible waiting, our threads can reduce load on the visible waiting array, and allow other threads and locks to benefit from using positive handover via the waiting array.
@@ 5 Related Work
CLH is extremely simple, has excellent Remote Memory Reference (RMR) complexity[stoc08-attiya, podc02-anderson], and requires just a single atomic exchange operation in the Acquire operation and no atomic read-modify-write instructions in Release. Unfortunately the waiting elements migrate between threads, which may be inimical to performance in NUMA environments. CLH locks also require explicit constructors and destructors, which may be inconvenient or preclude their use in some environments 555Many lock implementations require just trivial constructors to set the lock fields to 0 or some constant, and trivial destructors, which do nothing. The Linux kernel spinlock_t/qspinlock_t interface provides a constructor, but does not even expose a destructor. Such decisions influence and may limit which lock algorithms can be readily implemented under a particular interface.. Similarly, C++ std::mutex is allowed to be trivially destructible, meaning storage occupied by trivially destructible objects may be reused without calling the destructor. Under both GCC g++ version 15 and Clang++ Version 20, is_trivially_destructible reports True for std::mutex, and as such, destructors do not run.
@@
HemLock combines aspects of both CLH and MCS to form a lock that has very simple waiting node memory lifecycle constraints, is context-free but still scales well in common usage scenarios. HemLock requires just a singleton waiting element (queue node) per thread, which can be placed in thread-local storage. HemLock does not provide constant remote memory reference (RMR) complexity [opodis17-dvir] in scenarios where a thread holds multiple contended locks. In this situation the single node suffers from multi-waiting[spaa21-Dice]. Similar to MCS, HemLock lacks a constant-time unlock operation, whereas the unlock operator for CLH and Ticket locks is constant-time. Unlike MCS, HemLock requires active synchronous back-and-forth communication in the unlock path between the outgoing thread and its successor to protect the lifecycle of the waiting element. We note, however, that HemLock remains constant-time in the Release operator to the point where ownership is conveyed to the successor. HemLock uses address-based transfer of ownership, writing the address of the lock instead of a boolean, differentiating it from MCS and CLH. Reciprocating Locks[arxiv-recipro, ppopp25-dice], like HemLock, requires just a singleton per-thread waiting element allocated in thread-local storage. Reciprocating Locks has constant-time arrival and unlock paths, but is not FIFO, instead providing 2-bounded-bypass.
@@
Spinning : L indicates that waiters spin on a location that is usually collocated in the system topology with the waiting thread (typically on-stack or in thread-local storage) , which may confer performance benefits on some systems, such as those that use the Intel UPI[UPI] home-based coherence protocol, while U indicates that the location being waited upon is generally “unknown” and may be either local or remote to the waiting thread. P indicates private spinning, where at most one thread spins on a given location at a given time, G indicates global spinning, where all threads waiting on a given lock busy-wait on a single location, and S indicate semi-local[spaa21-Dice] spinning, where threads, absent hash collisions, enjoy private spinning.
@@
Context-free : indicates additional information does not need to be transferred from the Acquire operator to the corresponding Release operation.
@@
Explicit CTOR/DTOR Required : indicates the lock requires non-trivial constructors or destructors. CLH, when used in general purpose environments, for instance, requires destructors to run to release the wait elements referenced in the lock, to avoid memory leaks.
Thread CTOR/DTOR Required : indicates if per-thread constructors or destructors are needed to initialize or cleanup after dynamically created threads. We note that MCS implementations, could, if necessary, allocate queue nodes as needed, dynamically, in Acquire, and then release such queue nodes in Release. This avoids the need for thread-level destructors. (To avoid circular or recursive dependencies, we further require a specialized node allocator that is implemented via different types of locks). To avoid the allocator becoming a contention hotspot and bottleneck, common MCS implementations use per-thread caches of available elements, which need to be released when the thread terminates. Similarly, CLH could, in theory, allocate and free queue nodes on demand, although most implementations use a thread-local cache containing a singleton pointer to a free queue node, as seen in Scott’s [ScottB24] Figure 4.14. MCSH[MCSH] and K42[K42, ScottB24] MCS variations side step this concern by allocating elements on-stack, but at the cost of added complexity.
Max Remote Misses per episode : is the worst-case maximum number of misses to remote memory incurred, under simple sustained contention, by a matching Acquire-Release pair. Misses to remote memory – memory homed on a different NUMA node – may be more expensive than local misses on various platforms, such as modern Intel systems that use the UPI coherence fabric [UPI], where miss requests are first adjudicated by the home node of the cache line. Algorithms where nodes circulate are more vulnerable to accumulating such remote misses. For HemLock, we assume simple contention with no multi-waiting. We derived the “Max Remote Misses per episode” value via static inspection of the code, and determining which of the accesses that might cause coherence misses might also be to locations homed on remote nodes. We assume that back-to-back accesses of shared variables that reside on the same cache line, by a given thread, will result in just one miss. And for Hapax Locks we assume that the hapax value does not need to be replenished, which would incur an addition likely remote miss for the fetch-and-add atomic operation. We further assume no collisions (or near collisions, which could induce false sharing) in the waiting array.
@@
Dynamic Memory Allocation : indicates if the algorithm requires some type of dynamic memory allocation, typically for waiting elements. We note that MCS can avoid allocation and use on-stack allocation if all locking is lexically balanced and that CLH can also avoid allocation – by initially pairing elements 1:1 with locks and threads – if all threads and locks are statically allocated in advance of execution or where dynamically allocated threads and locks are effectively immortal. These advantageous circumstances are not normally found or available in general-purposes environments.
Trivial TryLock : indicates if the algorithm admits a trivial wait-free trylock operator. We note that Ticket-based locks can implement trylock if the Grant and Ticket fields are 64-bit bits, and effectively immune from roll-over overflow aliasing777Let’s say our ticket lock implementation has 32-bit Ticket and Grant fields. The naive but flawed way to implement trylock is to fetch the Grant field into a temporary , and then try to CAS Ticket from to . If execution is delayed between the load and the subsequent CAS, the values could have rolled around and we could suffer a false positive CAS success. This situation would be extremely rare, but not impossible. To avoid exclusion failure, if the CAS was successful, we might then try to adapt and remediate the algorithm to fetch and re-check Grant – after the successful CAS – and if Grant had changed, then we need to wait. This approach restricts us to 4G-1 possible concurrently waiting threads, but that’s arguably reasonable. But our trylock implementation is now flawed in the sense that it might need to wait, to recover. Another work-around for 32-bit fields is to collocate the Ticket and Grant fields within an atomic 64-bit superword and then to use a 64-bit CAS accordingly. This forces placement of the 2 fields, and can preclude attempts to limit false sharing by sequestering or isolating the constituent fields onto distinct cache lines. With 64-bit fields, however, given that rollover is not a practical concern, we can simply fetch Grant into and then try to CAS Ticket from to . The trylock attempt is successful if and only if the CAS was successful..